# End-to-End Autonomous Driving through V2X Cooperation

### v2

@@ Abstract
Cooperatively utilizing both ego-vehicle and infrastructure sensor data via V2X communication has emerged as a promising approach for advanced autonomous driving. However, current research mainly focuses on improving individual modules, rather than taking end-to-end learning to optimize final planning performance, resulting in underutilized data potential. In this paper, we introduce UniV2X, a pioneering cooperative autonomous driving framework that seamlessly integrates all key driving modules across diverse views into a unified network. We propose a sparse-dense hybrid data transmission and fusion mechanism for effective vehicle-infrastructure cooperation, offering three advantages: 1) Effective for simultaneously enhancing agent perception, online mapping, and occupancy prediction, ultimately improving planning performance. 2) Transmission-friendly for practical and limited communication conditions. 3) Reliable data fusion with interpretability of this hybrid data. We implement UniV2X, as well as reproducing several benchmark methods, on the challenging DAIR-V2X, the real-world cooperative driving dataset. Experimental results demonstrate the effectiveness of UniV2X in significantly enhancing planning performance, as well as all intermediate output performance. Code is at https://github.com/AIR-THU/UniV2X.
Keywords:
End-to-End Autonomous Driving V2X
[…]
1 Introduction
Table 1: Comparison with the existing methods for cooperative autonomous driving. "AgentP" denotes dynamic object perception. "Map" denotes online mapping. "Occ" denotes occupancy prediction. "-" denotes that the information is not provided or temporarily unable to verify. UniV2X targets planning output and conducts all essential tasks, and the transmitted data is effective, transmission-friendly, and reliable.
@@ 1 Introduction
V2VNet [47] Point Cloud BEV Feature - Medium No No
CoCa3D [18] Image BEV Feature - Medium No No
CoBEVT [51] Image BEV Feature - Medium No No
PP-VIC [63] Point Cloud Detected Result - High Yes No
DeepA [46] Image BEV Feature - Medium No No
TransIFF [6] Point Cloud Instance Feature - High Yes No
Where2Comm [17] Point Cloud Instance Feature - High Yes No
CSA [45] Image Raw Image Yes Low Yes Non-Explicit
COOPN [9] Point Cloud BEV Feature Yes Medium No Non-Explicit
@@
Figure 1: (a) VICAD: Infrastructure sensor installed highly has broad perception field [60, 63, 59, 58, 57, 14], which can supplement the blind and long-range spots of single vehicle. (b) Performance Enhancement: Compared with No Fusion solution, UniV2X achieves significant gains in various tasks, such as detection (+13%), mapping (+11.4%), occupancy prediction (+5.7%), and collision rate (-0.4%).
@@
Despite significant progress achieved through the integration of deep learning, single-vehicle autonomous driving still faces great safety challenges due to limited perceptual range and inadequate information, especially for vehicles relying on cost-effective cameras. Leveraging external sensors, particularly infrastructure sensors with a broader perception field, has shown promising potential for advancing autonomous driving capacities through Vehicle-to-Everything (V2X) communication. Several research studies have investigated the efficacy of external sensor data in diverse tasks such as detection [47, 18, 61, 44, 35], tracking [63], segmentation [51], localization [20, 10], and forecasting [63, 36, 40]. However, existing solutions primarily emphasize individual task optimization, neglecting the overall planning enhancement. This creates challenges in comprehensive data exploitation, driven by a misalignment between individual task goals and final planning objectives. Thus, end-to-end learning exploration, which directly optimizes the final planning output by harnessing both onboard and external sensor data, becomes necessary. In this paper, we focus on vehicle-infrastructure cooperative autonomous driving (VICAD), with images as input data. We also consider verifying UniV2X on V2V scenes.
@@
The VICAD problem can be formulated as a planning-centric optimization with multiple-view sensor inputs under constrained communication bandwidth. Compared with single-vehicle autonomous driving, VICAD poses additional challenges when addressed through end-to-end learning. Firstly, the transmitted infrastructure data must enhance both critical modules and the final planning performance in autonomous driving. These critical modules encompass dynamic obstacle perception, online mapping, and grid occupancy-based general obstacle detection, providing an explicit scene representation crucial for ensuring the safety of autonomous driving. Second, driven by real-time requirements and limited communication conditions, minimizing transmission costs becomes crucial to mitigate communication bandwidth consumption and reduce latency. Thirdly, transmitted data must be interpretable, allowing the vehicle to verify and judiciously employ the data to prevent potential safety issues such as communication attacks and data corruption. Addressing these challenges necessitates a well-designed solution for data transmission and cross-view data fusion.
@@
Here are a few straightforward attempts to address the cooperative driving problem through end-to-end learning. CSA [45] directly shares and feeds raw images received from other vehicles into basic neural networks for control output. CooperNaut [9] shares features derived from point clouds among vehicles and inputs them into a basic CNN network for the final output. However, these existing solutions rely on a vanilla approach, utilizing simple networks to optimize planning and control outputs. This paradigm lacks explicit modules, compromising safety assurance and interpretability. Especially within intricate urban settings, this approach falls short in ensuring the reliability of the driving system.
@@
To this end, we introduce UniV2X, an innovative cooperative autonomous driving framework that seamlessly integrates pivotal modules and cross views into a unified network, as depicted in Figure 2. Beyond the final planning task, we address three common tasks for scene representation in autonomous driving: 1) agent perception, encompassing 3D object detection, tracking, and motion forecasting for dynamic obstacle perception, 2) road element (especially lane) detection for online mapping, and 3) grid-occupancy prediction for general obstacle perception. Inspired by UniAD [16], we adopt a query-based architecture to establish connections across nodes, encompassing internal modules within infrastructure and ego-vehicle systems, as well as cross-view interactions. In transmission and cross-view interaction, we classify agent perception and road element detection as instance-level representation and occupancy prediction as scene-level representation. We transmit agent queries and lane queries for cross-view agent perception interaction and online mapping interaction. We transmit the occupied probability map, recognizing its dense nature at the scene-level occupancy, for cross-view occupancy interaction. This transmission, termed sparse-dense hybrid transmission, balances sparsity and density in spatial and feature dimensions, respectively. Cross-view data fusion, such as agent fusion, mainly involves temporal and spatial synchronization, cross-view data matching and fusion, data adaptation for planning and intermediate outputs. The resulting lightweight approach strengthens dynamic object perception, online mapping, and occupancy modules, thereby enhancing planning performance. Moreover, the interpretability of queries and occupied probability maps at the instance and scene levels, respectively, fortifies the reliability of the VICAD system, bolstering its transmission integrity and fusion safety.
@@
  • We reproduce several cooperative methods as benchmarks, as well as instantiating the UniV2X framework on DAIR-V2X[60] and V2X-Sim [25]. Experimental results underscore the efficacy of our end-to-end paradigm.
2 Related Works
2.0.1 Cooperative Autonomous Driving.
Leveraging V2X communication for cooperative autonomous driving has garnered significant attention. Works like [47, 26, 5, 54, 18, 31, 42] emphasize transmitting Bird’s Eye View (BEV) features extracted from point clouds or images to improve 3D object detection. [51] employs sparse transformers to enhance segmentation performance, while[36] focuses on utilizing sequential trajectories for motion forecasting. Addressing the latency issue in cooperation detection, [24, 62, 39, 56] utilize historical frames or features. Additionally, [17, 12, 6] transmit instance-level features or queries to reduce transmission costs and alleviate communication challenges. Datasets for cooperative object perception are provided by [60, 63, 52, 25, 33, 8, 53].  [63] releases a real-world trajectory dataset generated from infrastructure and vehicle sensor data. While most of these works concentrate on single tasks, a few also focus on end-to-end output planning or control tasks. For instance, [45] employs multi-vehicle images with a simple convolutional network to generate control outputs. Similarly,[9] adopts a simple MLP to learn planning outputs with BEV features from multi-vehicle point clouds. However, these solutions lack explicit interpretability modules. In this paper, we present UniV2X, a comprehensive framework that integrates all essential modules within a single model, utilizing end-to-end learning for optimizing final planning improvement as well as intermediate outputs.
2.0.2 End-to-End Autonomous Driving.
End-to-end autonomous driving involves the extraction of planning output directly from raw data in a differentiable and learnable manner. Pioneering work, such as [2], employs CNNs to generate control outputs from point cloud data. Others, exemplified by [37, 65], utilize point clouds and High-definition Maps as inputs. Concurrently, works like [38] leverage multi-modal sensor data as input to generate object density maps for visualization. For acquiring driving skills, [7, 34, 50] employ imitation learning (IL) to learn from expert demonstrations in an open-loop manner. In contrast, [30, 21, 19] utilize reinforcement learning (RL) to iteratively learn driving skills by interacting with the environment in a closed-loop fashion. Among related works, UniAD [16] is the first to use queries to connect all essential tasks such as perception, mapping, prediction, and planning. Through unifying these tasks into a single network and employing imitation learning, UniAD achieves remarkable performance on the nuScenes dataset [3]. However, UniAD solely considers single-vehicle sensor data, while our approach, UniV2X, leverages sensor data from diverse views.
3 Method
@@
Figure 2: Pipeline of Unified Autonomous Driving through V2X Cooperation (UniV2X). UniV2X aims to connect and jointly optimize all essential modules across diverse views for enhanced planning performance. Cross-view data interaction bolsters pivotal components in autonomous driving like agent perception, online mapping, and occupancy prediction. Additional flow prediction enables minimizing transmission cost for transmitting occupied probability map. Cross-view data fusion involves temporal and spatial synchronization, cross-view data matching and fusion, and data adaptation.
@@
In this section, we introduce UniV2X, all-in-one solution for vehicle and infrastructure cooperative autonomous driving (VICAD), together with the proposed sparse-dense hybrid data transmission and fusion design, as shown in Figure 2. We start by presenting the VICAD problem and introduce background. Following that, we describe how to generate sparse-dense hybrid data for transmission and cross-view data fusion. The training process is also outlined in this section.
3.1 VICAD Problem Formulation
@@
The VICAD problem is planning-oriented, aiming to improve the planning performance by utilizing both infrastructure sensor data and ego-vehicle sensor data thorough V2X communication. This paper focuses on the images as inputs. The input of VICAD consists of two parts: (a) Ego-vehicle images $\{I_{v}(t)|t\leq t_{v}\}$ and the relative pose $M_{v}(t_{v})$ at the current vehicle timestamp $t_{v}$. (b) Infrastructure images $\{I_{i}(t)|t\leq t_{i}\}$ and the relative pose $M_{i}(t_{i})$ at the current infrastructure timestamp $t_{i}$. Note that in practical scenario, the timestamp $t_{i}$ should be earlier than timestamp $t_{v}$ due to the communication latency. The output of VICAD is to predict future coordinates of ego vehicle for time steps $t=t_{v}+1,\cdots,t_{pred}$.
3.1.1 Evaluation Metrics.
@@
We evaluate the planning performance with L2 Error and Collision Rate, and measure transmission cost with Bytes Per Second (BPS), as suggested in [60, 63]. We provide the detailed explanations in the Appendix.
3.1.2 Challenges.
@@
Compared with single-vehicle autonomous driving, VICAD presents additional challenges: (1) Limited by practical communication conditions, fewer infrastructure data should be transmitted to vehicle to minimize bandwidth usage and reduce latency. (2) Wireless communication introduce unavoidable latency, potentially leading to temporal misalignment in data fusion. (3) Potential communication attacks and data corruption can render transmitted data untrustworthy. This highlights the need for interpretable transmitted data.
3.2 Background
3.2.1 Explicit Scene Representation.
Planning and decision-making built on explicit scene representation facilitates explainable and trustworthy autonomous driving. Here are several common and vital modules in autonomous driving.
  • Agent perception is used for accurately sensing dynamic obstacles (3D Object Detection), correlating obstacles (Tracking), and ultimately predicting their future behavior (Motion Forecasting). Accurate sensing and prediction help avoid collisions and ensure driving safety.
  • Online mapping aims at extracting road elements to construct a map in real time, which can be used for positioning and making decisions more compliant with traffic rules like Tesla way [41]. It is considered to be an effective alternative to HD (high-precision) maps, especially for low-cost solutions.
  • Grid-based occupancy prediction is used for general obstacle detection by predicting the occupancy of each grid [48, 43]. This is regarded as the effective solution for long-tailed challenges in autonomous driving.
3.2.2 Data for Transmission.
It involves three primary types in V2X cooperation:
  • Raw data like raw images. This data type maintains all valuable information but requires significant transmission costs.
  • Perception outputs like detection results. While communication-friendly, this data is not suitable for seamless integration into an end-to-end framework.
  • Intermediate-level data such as Bird’s Eye View (BEV) features and queries [12, 6]. Compared to the above two data, this data achieves a balance between preserving valuable information and reducing redundant transmission.
@@ 3.2 Background
To ensure effective, transmission-friendly, and reliable transmitted data, we propose a sparse-dense hybrid transmission mechanism. Queries, as lightweight instance-level features, enhance agent perception and online mapping, as dynamic obstacles and lanes can be treated as instance-level representations. Occupied probability maps, channel-sparse scene-level features, improve occupancy prediction. Compared to less interpretable and high-cost BEV features, occupied probability maps offer pixel-level interpretability and lower transmission costs.
3.3 Sparse-Dense Hybrid Data Generation
@@
This part illustrates how to generate sparse-dense hybrid data for transmission in infrastructure system. As depicted in Figure 2, the infrastructure system is composed of: Backbone, TrackFormer, MapFormer, OccFormer, and OccFlowNet.
@@
BEVFormer[27] is adopted as the backbone to extract image features and transform them into bird’s-eye-view (BEV) features $B_{inf}$ with size of (200, 200, 256) by incorporating spatial cross-attention and temporal self-attention. TrackFormer is based on DETR [4], which optimizes detection and multi-object tracking together, eliminating the need for non-differentiable post-processing like NMS [1]. The ultimate filtered output from TrackFormer contains $N_{a}^{inf}$ valid agent queries $\{Q_{A}^{inf}\}$ with a feature dimension of 256 and their corresponding assigned tracking IDs and reference points. MapFormer is based on Panoptic SegFormer [28]. We mainly focus on the lane line and crosswalk elements. During transmission, we filter out low-scoring queries using boxes generated from the classification decoder, and exclusively transmit $N_{l}^{inf}$ valid lane queries $\{Q_{L}^{inf}\}$ with a feature dimension of 256, along with their corresponding reference points. Original OccFormer in UniAD [16] solely considers instance-level occupancy associated with agent queries, predicting multiple steps. However, occupancy serves as a complementary factor to object perception for general obstacle detection, and transmitting multiple probability maps incurs significant transmission costs. To address these challenges, we retain the dense feature obtained through pixel-level attention with a size of (200, 200, 256). Initially, a Multi-layer Perception (MLP) is employed to transform the dense feature into BEV occupied probability map denoted as $P^{inf}$ with a size of (200, 200). Subsequently, adopting the feature flow prediction approach [62, 63], an additional probability flow module is utilized to represent T-step maps via a linear operation as
@@
3.3.1 Advantages of Sparse-Dense Hybrid Data Transmission.
Agent queries and lane queries, as instance-level features, exhibit remarkable sparsity in the spatial dimension. Conversely, the BEV occupied probability map is dense in the spatial dimension but is characterized by a single dimension in terms of feature dimension. The probability flow mirrors the sparse-dense characteristics of their respective counterparts, the probability map. Hence, the sparse-dense hybrid data transmission incurs minimal transmission costs and is communication-friendly. A transmission cost comparison is also presented in Table 2.
Table 2: Transmission Cost in Each Transmission. Compared with raw images and BEV feature, hybrid data transmission requires much less transmission cost.
Data Type Size Transmission Cost (Bytes)
Raw Image (1080, 1920, 3) $\sim 2\times 10^{8}$
BEV Feature (200, 200, 256) $\sim 4\times 10^{8}$
Hybrid Feature ($N_{a}^{inf}$ + $N_{l}^{inf}$, 256) $\times$ 2, (200, 200) $\times$ 2 $\sim 10^{5}$
3.4 Cross-View Data Fusion (Agent Fusion)
In the ego-vehicle system, the BEV features $B_{veh}$ are first extracted from the images captured by onboard sensors. We also adopt TrackFormer, MapFormer, and OccFormer to generate the corresponding agent queries $\{Q_{A}^{veh}\}$, lane queries $\{Q_{L}^{veh}\}$, and the occupied probability map $P^{veh}$. The network for these modules aligns with that of the infrastructure system. In this section, we describe how to implement cross-view agent fusion. Cross-view agent fusion is mainly composed of temporal synchronization for latency compensation, spatial synchronization to unify the cross-view coordinates, data matching and fusing, and data adaptation for planning and intermediate outputs. Some modules are similar in implementing lane fusion and occupancy fusion.
3.4.1 Temporal Synchronization with Flow Prediction.
The transmission delay in wireless communication, as $t_{i}$ is earlier than $t_{v}$, is significant in complex traffic systems, especially for the busy intersection scenario. Due to the movement of dynamic objects, there is a temporal misalignment when fusing data from different sources. To address that, we incorporate feature prediction into infrastructure agent queries to mitigate latency, following feature flow prediction as [62, 63]. Specifically, we input both agent query $Q_{A}^{inf}$ and query associated in the previous frame into QueryFlowNet, a three-layer Multi-Layer Perceptron (MLP), to generate the agent query flow $Q_{AFlow}^{inf}$. The dimensions of the agent query flow match those of the agent query. Subsequently, a linear operation forecasts future features can be used to mitigate latency $t_{v}-t_{i}$, depicted as
@@ 3.4 Cross-View Data Fusion (Agent Fusion)
Notably, QueryFlowNet of the Flow Prediction module is not trained in an end-to-end manner in UniV2X. We adopt self-supervised learning following [62, 63].
3.4.2 Spatial Synchronization with Rotation-Aware Query Transformation.
@@
3.4.3 Cross-View Query Matching and Fusion.
At this stage, cross-view agent queries are temporally and spatially synchronized. To match corresponding queries from different sides, we calculate the Euclidean distance of their reference points and employ the Hungarian method [22]. For the matched query pairs $Q_{A}^{inf}$ and $Q_{A}^{veh}$, they are fed into a three-layer MLP to generate the cooperated query $Q_{A}$, which is used to update the ego-vehicle agent query $Q_{A}^{veh}$. For the unmatched queries from infrastructure, they are utilized to be added to the ego-vehicle queries. Finally, we assign tracking IDs and filter out cross-view fused queries with low detection confidence, resulting in the ultimate agent queries.
3.4.4 Ego Identification and Removing.
This module is used to eliminate the problem of false detection in the ego-vehicle area. From the infrastructure view, the ego vehicle can be perceived either as a distinct obstacle in agent perception or as part of the occupied area in occupancy prediction. Following cross-view data fusion, there is a possibility of generating an obstacle query within the region where the ego vehicle is located, thereby marking the ego-vehicle area as occupied. Such an occurrence can significantly disrupt decision-making processes and ultimately impact decision-making performance. To mitigate this issue, we define the ego-vehicle area as a rectangle, filter queries within this area, and designate this region as unoccupied. This configuration proves effective in enhancing planning performance, particularly in the L2 Error evaluation. However, this straightforward solution may not consistently perform optimally due to relative position errors between infrastructure and ego vehicle, stemming from positioning and calibration inaccuracies [55, 13]. Further exploration and refinement are essential to contribute to cooperative autonomous driving.
3.4.5 Decoder Input Augmentation for Intermediate Output.
@@
3.5 Cross-View Data Fusion (Lane Fusion)
@@
3.6 Cross-View Data Fusion (Occupancy Fusion)

### v3

@@ Abstract
Cooperatively utilizing both ego-vehicle and infrastructure sensor data via V2X communication has emerged as a promising approach for advanced autonomous driving. However, current research mainly focuses on improving individual modules, rather than taking end-to-end learning to optimize final planning performance, resulting in underutilized data potential. In this paper, we introduce UniV2X, a pioneering cooperative autonomous driving framework that seamlessly integrates all key driving modules across diverse views into a unified network. We propose a sparse-dense hybrid data transmission and fusion mechanism for effective vehicle-infrastructure cooperation, offering three advantages: 1) Effective for simultaneously enhancing agent perception, online mapping, and occupancy prediction, ultimately improving planning performance. 2) Transmission-friendly for practical and limited communication conditions. 3) Reliable data fusion with interpretability of this hybrid data. We implement UniV2X, as well as reproducing several benchmark methods, on the challenging DAIR-V2X, the real-world cooperative driving dataset. Experimental results demonstrate the effectiveness of UniV2X in significantly enhancing planning performance, as well as all intermediate output performance. The project is available at https://github.com/AIR-THU/UniV2X.
Introduction
Table 1: Comparison with the existing methods for cooperative autonomous driving. “AgentP” denotes dynamic object perception.“Map” denotes online mapping. “Occ” denotes occupancy prediction. “-” denotes that the information is not verified.
@@ Introduction
V2VNet (Wang et al. 2020) Point Cloud BEV Feature - Medium No No
CoCa3D (Hu et al. 2023a) Image BEV Feature - Medium No No
CoBEVT (Xu et al. 2022a) Image BEV Feature - Medium No No
PP-VIC (Yu et al. 2023c) Point Cloud Detected Result - High Yes No
DeepA (Wang et al. 2023) Image BEV Feature - Medium No No
TransIFF (Chen, Shi, and Jia 2023) Point Cloud Instance Feature - High Yes No
Where2Comm (Hu et al. 2022) Point Cloud Instance Feature - High Yes No
CSA (Valiente et al. 2019) Image Raw Image Yes Low Yes Non-Explicit
CooperNaut (Cui et al. 2022) Point Cloud BEV Feature Yes Medium No Non-Explicit
@@
Despite significant progress achieved through the integration of deep learning, single-vehicle autonomous driving still faces great safety challenges due to limited perceptual range and inadequate information, especially for vehicles relying on cost-effective cameras. Leveraging external sensors, particularly infrastructure sensors with a broader perception field, has shown promising potential for advancing autonomous driving capacities through Vehicle-to-Everything (V2X) communication (see Fig. 1 (a)). Several research studies have investigated the efficacy of external sensor data in diverse tasks such as detection (Wang et al. 2020; Hu et al. 2023a; Yu et al. 2023a; Tianhang et al. 2023; Qiu et al. 2022), tracking (Yu et al. 2023c), segmentation (Xu et al. 2022a), localization (Jiang et al. 2023; Dong et al. 2023), and forecasting (Yu et al. 2023c; Ruan et al. 2023; Song et al. 2024). However, existing solutions primarily emphasize individual task optimization, neglecting the overall planning enhancement. This creates challenges in comprehensive data exploitation, driven by a misalignment between individual task goals and final planning objectives. Thus, end-to-end learning exploration, which directly optimizes the final planning output by harnessing both onboard and external sensor data, becomes necessary. In this paper, we focus on vehicle-infrastructure cooperative autonomous driving (VICAD).
Figure 1: (a) VICAD: Infrastructure sensor installed highly has a broad perception field (Yu et al. 2022, 2023c; Yang et al. 2023d), which can supplement the blind and long-range spots of single vehicle. (b) Performance Enhancement: Compared with No Fusion solution, UniV2X achieves significant gains in various tasks, such as detection (+13%), mapping (+11.4%), occupancy prediction (+5.7%), and collision rate (-0.5%).
@@
The VICAD problem can be formulated as a planning-centric optimization with multiple-view sensor inputs under constrained communication bandwidth. Compared with single-vehicle autonomous driving, VICAD poses additional challenges when addressed through end-to-end learning. Firstly, the transmitted infrastructure data must be effective. It should enhance both critical modules and the final planning performance in autonomous driving. These critical modules encompass dynamic obstacle perception, online mapping, and grid occupancy-based general obstacle detection, providing an explicit scene representation crucial for ensuring the safety of autonomous driving. Second, the data must be friendly. Driven by real-time requirements and limited communication conditions, minimizing transmission costs becomes crucial to mitigate communication bandwidth consumption and reduce latency. Thirdly, the transmitted data must be reliable. it should be reliable. Vehicles need interpretable information that can be validated and used judiciously to avoid safety issues such as communication attacks or data corruption. Addressing these challenges necessitates a well-designed solution for data transmission and cross-view data fusion.
@@
Here are a few straightforward attempts to address the cooperative driving problem through end-to-end learning. CSA (Valiente et al. 2019) directly shares and feeds raw images received from other vehicles into basic neural networks for control output. CooperNaut (Cui et al. 2022) shares features derived from point clouds among vehicles and inputs them into a basic CNN network for the final output. However, these existing solutions rely on a vanilla approach, utilizing simple networks to optimize planning and control outputs. This paradigm lacks explicit modules, compromising safety assurance and interpretability. Especially within intricate urban settings, this approach falls short in ensuring the reliability of the driving system. More comparisons are in Table 1, and related work is discussed in the appendix.
@@
To this end, we introduce UniV2X, an innovative cooperative autonomous driving framework that seamlessly integrates pivotal modules and cross views into a unified network, as depicted in Figure 2. Beyond the final planning task, we address three common tasks for scene representation in autonomous driving: 1) agent perception, encompassing 3D object detection, tracking, and motion forecasting for dynamic obstacle perception, 2) road element (especially lane) detection for online mapping, and 3) grid-occupancy prediction for general obstacle perception. Inspired by UniAD (Hu et al. 2023b), we adopt a query-based architecture to establish connections across nodes, encompassing internal modules within infrastructure and ego-vehicle systems, as well as cross-view interactions. In transmission and cross-view interaction, we classify agent perception and road element detection as instance-level representation and occupancy prediction as scene-level representation. We transmit agent queries and lane queries for cross-view agent perception interaction and online mapping interaction. We transmit the occupied probability map, recognizing its dense nature at the scene-level occupancy, for cross-view occupancy interaction. This transmission, termed sparse-dense hybrid transmission, balances sparsity and density in spatial and feature dimensions, respectively. Cross-view data fusion, such as agent fusion, mainly involves temporal and spatial synchronization, cross-view data matching and fusion, data adaptation for planning and intermediate outputs. The resulting lightweight approach strengthens dynamic object perception, online mapping, and occupancy modules, thereby enhancing planning performance. Moreover, the interpretability of queries and occupied probability maps at the instance and scene levels, respectively, fortifies the reliability of the VICAD system, bolstering its transmission integrity and fusion safety.
@@
  • We reproduce several cooperative methods as benchmarks, as well as instantiating the UniV2X on DAIR-V2X (Yu et al. 2022). Experimental results underscore the efficacy of our end-to-end paradigm (see Fig. 1 (b)).
Method
@@
Figure 2: Pipeline of Unified Autonomous Driving through V2X Cooperation (UniV2X). UniV2X aims to connect and jointly optimize all essential modules across diverse views for enhanced planning performance. Cross-view data interaction bolsters pivotal components in autonomous driving like agent perception, online mapping, and occupancy prediction. Additional flow prediction enables minimizing transmission costs for transmitting occupied probability map. Cross-view data fusion involves temporal and spatial synchronization, cross-view data matching and fusion, and data adaptation.
@@
In this section, we introduce UniV2X, an all-in-one solution for vehicle and infrastructure cooperative autonomous driving (VICAD), together with the proposed sparse-dense hybrid data transmission and fusion design, as shown in Figure 2. We start by presenting the VICAD problem and introduce the background. Following that, we describe how to generate sparse-dense hybrid data for transmission and cross-view data fusion. The training process is also outlined.
VICAD Problem Formulation
@@
The VICAD problem is planning-oriented, aiming to improve planning performance by utilizing both infrastructure sensor data and ego-vehicle sensor data through V2X communication. This paper focuses on the images as inputs. The input of VICAD consists of two parts: (a) Ego-vehicle images $\{I_{v}(t)|t\leq t_{v}\}$ and the relative pose $M_{v}(t_{v})$ at the current vehicle timestamp $t_{v}$. (b) Infrastructure images $\{I_{i}(t)|t\leq t_{i}\}$ and the relative pose $M_{i}(t_{i})$ at the current infrastructure timestamp $t_{i}$. Note that in a practical scenario, the timestamp $t_{i}$ should be earlier than the timestamp $t_{v}$ due to the communication latency. The output of VICAD is to predict future coordinates of ego vehicle for time steps $t=t_{v}+1,\cdots,t_{pred}$.
Evaluation Metrics.
@@
We evaluate the planning performance with L2 Error, Collision Rate and Off-Road Rate, and measure transmission cost with Bytes Per Second (BPS), as suggested in (Yu et al. 2022, 2023c). We provide detailed explanations for these metrics in the Appendix.
Challenges.
@@
Compared with single-vehicle autonomous driving, VICAD presents additional challenges: (1) Limited by practical communication conditions, fewer infrastructure data should be transmitted to vehicles to minimize bandwidth usage and reduce latency. (2) Wireless communication causes latency, potentially leading to temporal misalignment in data fusion. (3) Potential communication attacks and data corruption can render transmitted data untrustworthy. This highlights the need for interpretable transmitted data.
Data for Transmission.
@@
It involves three primary types in V2X cooperation: raw data like raw images, perception outputs like detection results, and intermediate-level data such as Bird’s Eye View (BEV) features and queries (Fan et al. 2024; Chen, Shi, and Jia 2023; Zhong et al. 2024). Compared to raw data and detection results, intermediate-level data achieves a balance between preserving valuable information and reducing redundant transmission. To ensure effective, transmission-friendly, and reliable transmitted data, we propose a sparse-dense hybrid transmission mechanism. Queries, as lightweight instance-level features, enhance agent perception and online mapping, as dynamic obstacles and lanes can be treated as instance-level representations. Occupied probability maps, channel-sparse scene-level features, improve occupancy prediction. Compared to less interpretable and high-cost BEV features, occupied probability maps offer pixel-level interpretability and lower transmission costs.
Sparse-Dense Hybrid Data Generation
@@
This part illustrates how to generate sparse-dense hybrid data for transmission in infrastructure system.
@@
BEVFormer(Li et al. 2022b) is adopted as the backbone to extract image features and transform them into bird’s-eye-view (BEV) features $B_{inf}$ with size of (200, 200, 256) by incorporating spatial cross-attention and temporal self-attention. TrackFormer is based on DETR (Carion et al. 2020), which optimizes detection and multi-object tracking together, eliminating the need for non-differentiable post-processing like NMS (Bodla et al. 2017). The ultimate filtered output from TrackFormer contains $N_{a}^{inf}$ valid agent queries $\{Q_{A}^{inf}\}$ with a feature dimension of 256 and their corresponding assigned tracking IDs and reference points. MapFormer is based on Panoptic SegFormer (Li et al. 2022c). We mainly focus on the lane line and crosswalk elements. During transmission, we filter out low-scoring queries using boxes generated from the classification decoder, and exclusively transmit $N_{l}^{inf}$ valid lane queries $\{Q_{L}^{inf}\}$ with a feature dimension of 256, along with their corresponding reference points. Original OccFormer in UniAD (Hu et al. 2023b) solely considers instance-level occupancy associated with agent queries, predicting multiple steps. However, occupancy serves as a complementary factor to object perception for general obstacle detection, and transmitting multiple probability maps incurs significant transmission costs. To address these challenges, we retain the dense feature obtained through pixel-level attention with a size of (200, 200, 256). Initially, a Multi-layer Perception (MLP) is employed to transform the dense feature into BEV occupied probability map denoted as $P^{inf}$ with a size of (200, 200). Subsequently, adopting the feature flow prediction approach (Yu et al. 2023b, c), an additional probability flow module is utilized to represent T-step maps via a linear operation as
@@
Cross-View Data Fusion (Agent Fusion)
In the ego-vehicle system, the BEV features $B_{veh}$ are first extracted from the images captured by onboard sensors. We also adopt TrackFormer, MapFormer, and OccFormer to generate the corresponding agent queries $\{Q_{A}^{veh}\}$, lane queries $\{Q_{L}^{veh}\}$, and the occupied probability map $P^{veh}$. The network for these modules aligns with that of the infrastructure system. In this section, we describe how to implement cross-view agent fusion. Cross-view agent fusion is mainly composed of temporal synchronization for latency compensation, spatial synchronization to unify the cross-view coordinates, data matching and fusing, and data adaptation for planning and intermediate outputs.
Temporal Synchronization with Flow Prediction.
The transmission delay in wireless communication, as $t_{i}$ is earlier than $t_{v}$, is significant in complex traffic systems, especially for the busy intersection scenario. Due to the movement of dynamic objects, there is a temporal misalignment when fusing data from different sources. To address that, we incorporate feature prediction into infrastructure agent queries to mitigate latency, following feature flow prediction as (Yu et al. 2023b, c). Specifically, we input both agent query $Q_{A}^{inf}$ and query associated in the previous frame into QueryFlowNet, a three-layer Multi-Layer Perceptron (MLP), to generate the agent query flow $Q_{AFlow}^{inf}$. The dimensions of the agent query flow match those of the agent query. Subsequently, a linear operation forecasts future features can be used to mitigate latency $t_{v}-t_{i}$, depicted as
@@
Notably, QueryFlowNet of the Flow Prediction module is not trained in an end-to-end manner in UniV2X. We adopt self-supervised learning following (Yu et al. 2023b, c).
Spatial Synchronization with Rotation-Aware Query Transformation.
@@
Cross-View Query Matching and Fusion.
At this stage, cross-view agent queries are temporally and spatially synchronized. To match corresponding queries from different sides, we calculate the Euclidean distance of their reference points and employ the Hungarian method (Kuhn 1955). For the matched query pairs $Q_{A}^{inf}$ and $Q_{A}^{veh}$, they are fed into a three-layer MLP to generate the cooperated query $Q_{A}$, which is used to update the ego-vehicle agent query $Q_{A}^{veh}$. For the unmatched queries from infrastructure, they are utilized to be added to the ego-vehicle queries. Finally, we assign tracking IDs and filter out cross-view fused queries with low detection confidence, obtaining the final agent queries.
Ego Identification and Removing.
This module is used to eliminate the problem of false detection in the ego-vehicle area. From the infrastructure view, the ego vehicle can be perceived either as a distinct obstacle in agent perception or as part of the occupied area in occupancy prediction. Following cross-view data fusion, there is a possibility of generating an obstacle query within the region where the ego vehicle is located, thereby marking the ego-vehicle area as occupied. Such an occurrence can significantly disrupt decision-making processes and ultimately impact decision-making performance. To mitigate this issue, we define the ego-vehicle area as a rectangle, filter queries within this area, and designate this region as unoccupied. However, this straightforward solution may not consistently perform optimally due to relative position errors between infrastructure and ego vehicle, stemming from positioning and calibration inaccuracies (Yang et al. 2023b; Gu et al. 2023). Further exploration and refinement are essential to contribute to cooperative autonomous driving.
Decoder Input Augmentation for Intermediate Output.
@@
Cross-View Data Fusion (Lane Fusion)
@@
Cross-View Data Fusion (Occupancy Fusion)